a grocer mixes 5 pound of egg noddels costing $.80 per pound with 2 pounds of spinach noodles costing $1.50 per pound. What is the cost per pound of the mixture.
Answer by josmiceli(19441)

5 lb x .80/lb = $4.00
this is the cost of 5 lb of egg noodles
2 lb x 1.50/lb = $3.00
this is the cost of 2 lb of spinich noodles
there is 7 lb of the mixture
($4.00 + $3.00)/ 7 lb = cost per pound of the mixture
$1.00/ lb = cost per pound of the mixture